the bank loaned out $17,000, part of it at the rate of 8% per year and the rest at 18% per year. If the interest received in one year totaled $2000, how much was loaned at 8%?
How much of the $17,000 did the bank loan out at 8%?

Answer by stanbon(75887)

the bank loaned out $17,000, part of it at the rate of 8% per year and the rest at 18% per year. If the interest received in one year totaled $2000, how much was loaned at 8%?
How much of the $17,000 did the bank loan out at 8%?
-------
Quantity Eq::: x + y = 17000
Interest Eq:::0.08x + 0.18y = 2000
----------------------------------
Multiply thru 1st by 8
Multiply thru 2nd by 100
------
8x + 8y = 8*17000
8x +18y = 100*2000
------
10y = 64000
y = $6400 (amt. invested at 18%
----
Solve for "x" using x + y = 17000
x = 17000-6400 = $10,600 (amt. invested at 8%)
------
Cheers,
Stan H.
=================